Hi. I need some assistance on how to work this problem out.
Consider this system:
{3(x+y)=x-3
-y=2x+3/3
I need to figure out if they are dependent or independent and if it's consistent or inconsistent. Thank you for your time.

Answer by ewatrrr(10682)

 
Hi,
A consistent system has at least one solution.
An inconsistent system has no solutions.
An independent system has exactly one solution.
A dependent system has infinitely many solutions.
3(x+y)=x-3
-y = (2x+3)/3
2x+3y = -3
-2x-3y = 3
0 = 0 Dependent system, has infinitely many solutions
Graphically these EQs represent same Line